OpenStudy (anonymous):

1/2(20 x 5)+200 x 5 = P

OpenStudy (anonymous):

210x^5

OpenStudy (anonymous):

1050?

OpenStudy (anonymous):

'x' is multiplication sign or variable?

OpenStudy (anonymous):

Muiltiplication

OpenStudy (anonymous):

it will b 1050...if its the sign

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }=.5\] So, \[.5(20\times5)+200 \times 5=P\] Use distributive property of multiplication to solve .5(20 * 5); \[.5 \times 20 = 10\] \[.5 \times 5 = 2.5\] so, \[10 \times 2.5 + 200 \times 5 = P\] \[10 \times 2.5 = 25\] and \[200 \times 5 = 1000\] So, \[25+1,000=?\] And that's your answer.

OpenStudy (anonymous):

I see. Use * next time. And of course, it is 1050. Do you have any question or need something to cleared?

OpenStudy (anonymous):

No, thanks for your time people.

OpenStudy (anonymous):

@micahwood50 @pkjha3105 How is it 1050? It seems as thought you forgot to use distributing of .5 and 5 rather and just multiply 10 and 5 to get the 50.

OpenStudy (anonymous):

@I4NI its 50+1000=1050

OpenStudy (anonymous):

@I4NI Have you ever heard of order of operations? You have to solve something in parentheses first.

OpenStudy (anonymous):

i thnk .5(100)=50

OpenStudy (anonymous):

Ahh, now I see, thanks for the correction. And to the OP it is 1050.

OpenStudy (anonymous):

i hnnt multiplied by frction ,that i ll need to consider abt othr term

OpenStudy (anonymous):

Ok, thanks again..